OpenStudy (anonymous):

Annual payments with 5% interest compounded annually for 8 years to accumulate $600

OpenStudy (anonymous):

600=x(1+.05)^8 then x=600/1.05^8=406

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

ur welcome